OpenStudy (anonymous):

How many grams of carbon dioxide can form when a mixture of 3.60 g ethylene (C2H4) and 4.10 g of oxygen is ignited, assuming complete combustion to form carbon dioxide and water?

OpenStudy (aaronq):

Stoichiometry When performing stoichiometric calculations the values to be used \(\sf \color{red}{must~be~in~moles}\). This is because the Molar Masses (\(\sf M_{m}\) of two elements are not the same. This is to say that 1 gram of Hydrogen atoms is not the same number of atoms as 1 gram of Nitrogen atoms. We need to be working with numbers of particles (e.g. atoms, molecules, formula units) and using moles is the way this is done. --------------------------------------------------- To interconvert between mass and moles, use the relationship: \(\Large \sf moles=\dfrac{mass}{Molar~mass}\) --------------------------------------------------- General Scheme for Stoichiometric Calculations: \(\sf \large 1.\)First write and balance an equation for the process described. \(\sf \large 2.\)Next, use the stoichiometric coefficients to find moles produced. \(\sf \large 3.\) Set up a ratio using the species of interest, like so (for a general reaction): \(\sf \large \color{red}{a}A + \color{blue}{b}B \rightleftharpoons \color{green}{c}C\) here upper case letters (A,B,C) are the chemical species, and lower case letters (a,b,c) are the coefficients, \(\sf \dfrac{moles_A}{\color{red}{a}}=\dfrac{moles_B}{\color{blue}{b}}=\dfrac{moles _C}{\color{green}{c}}\) You can use any pair from here, depending on what the question is asking for. For example: if you have 2 moles of B, how many moles of C can you produce? Now rearrange algebraically, \(\sf\dfrac{2}{\color{blue}{b}}=\dfrac{moles_C}{\color{green}{c}}\rightarrow moles _C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) \(\sf \large 4.\) Solve

OpenStudy (welshfella):

so first you need to balance the equation nC2H4 + mO2 --> pCO2 + qH2O HinT:- start with n = 2 and p = 4